OpenStudy (anonymous):

Caden and Madison share a reward of $180 in a ratio of 7 : 2 what fraction does Madison get

jimthompson5910 (jim_thompson5910):

"Caden and Madison share a reward of $180 in a ratio of 7 : 2" means that we have this equation 7x+2x = 180 are you able to solve for x here?

OpenStudy (anonymous):

20

OpenStudy (anonymous):

tanks you

OpenStudy (anonymous):

thank*

jimthompson5910 (jim_thompson5910):

madison's share is 2x, so 2*x = 2*20 = 40 is the amount of money she gets